OpenStudy (anonymous):

The length of a rectangle is 4 m less than three times the width, and the area of the rectangle is 55 m2 . Find the dimensions of the rectangle.

OpenStudy (merchandize):

According to this statement The length of a rectangle is 4 m less than three times the width, and the area of the rectangle is 55 m2 we have to make an equation

OpenStudy (anonymous):

so how should i do it

OpenStudy (merchandize):

read it carefully and the equation will be (3w-4)w = 55m^2 Solve the equation and get the value of W and L

OpenStudy (anonymous):

i have no idea how do solve that equation

OpenStudy (anonymous):

do i solve for w?

OpenStudy (merchandize):

A=W*L, L=A/W this way you can find the value of L

OpenStudy (merchandize):

yes solve it for W first

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i think i need more help

OpenStudy (anonymous):

with this

OpenStudy (merchandize):

Ask your doubt

OpenStudy (mertsj):